Fermi National Laboratory
 

Operation Summary for May 13, 2011 - May 16, 2011

The Integrated Luminosity for the period from 5/9/11 to 5/16/11 was 63.88 pb-1 (inverse picobarns). NuMI took 4.15E18 of beam during this same time period.

Friday May 13

The day shift began with Operations monitoring the Tevatron (TEV) store 8733 (50.9E30) and the Recycler stash (180.5E10), with the Antiproton Source (Pbar) stacking (11.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, TEV personnel conducted Tevatron Electron Lens (TEL2) hollow beam studies from 11 AM to 2:30 PM. Operations terminated store 8733 at 2:30 PM and put the TEV into shot setup. Operations reported some downtime due to a Booster West Anode Power Supply trip; it reset.

During the evening shift, Operations established store 8734 at 4:10 PM with an average initial luminosity (AIL) of 312.58E30. Operators reported problems turning on TEL1 at the beginning of the store; EE Support personnel assisted Operations in getting this device on. Operations disabled beam to T-979 at 6:23 PM. Operators reported that the Booster low-level RF (LLRF) room temperature came into alarm; operators cycled the disconnect for the air conditioning (AC) unit to get the compressors firing again.

Saturday May 14

The midnight shift began with Operations monitoring the store 8734 (105.54E30) and the stash (221.9E10), with Pbar stacking (17.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8734 at 7:26 AM and put the TEV into shot setup. Operators reported that the Booster LLRF room temperature remained problematic all shift, requiring attention. Operators reported filling the E48 kicker with SF6 gas.

During the day shift, Operations established store 8735 at 8:43 AM with an AIL of 362.09E30. Operators rebooted the TEV FBI and the SWIC front ends multiple times during the shift. Operators reported that the E48 kicker SF6 leak rate had increased; kicker personnel determined that repairs would be needed on Monday, May 16, and that operators should top off the kicker with SF6 before each shot setup. Operations reported continuing problems with the Booster LLRF room temperature.

During the evening shift, Operations terminated T-979 beam at 6 PM. Operators reported that the TEV orbits from the previous store looked a bit odd; TEV personnel checked things out and recommended no orbit corrections for the next store. Operations terminated store 8735 at 11:29 PM and put the TEV into shot setup.

Sunday May 15

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash, and the Pbar stack, and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators reported topping off the E48 kicker with SF6. Operations established store 8736 at 12:44 AM with an AIL of 342.4E30. The Booster LLRF room temperature continued to cause problems. Operators reported that the MI SBD died just before shot setup and did not come back on cleanly.

During the day shift, Operations established beam to T-979 at 8:17 AM. Operations reported that the FBI required multiple reboots throughout the shift. The H- Source high voltage tripped off several times during the shift, but did not cause any downtime due to the fact that the I- Source was the active source.

During the evening shift, Operations terminated store 8736 at 4:10 PM, put the TEV into shot setup, and established store 8737 at 5:33 PM with an AIL of 381.38E30. Operations terminated T-979 beam at 6 PM. Controls personnel looked into and fixed the problems with the MI SBD. The H- Source high voltage tripped off. FESS and Booster personnel looked at the Booster LLRF room AC unit and made adjustments. Operations reported that the TEV FBI only needed one reboot this shift.

Monday May 16

The midnight shift began with Operations monitoring store 8737 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators reported that various Booster and MI RF station trips caused some downtime. Operations reported that the analog to digital (A/D) chassis for some of the Linac quadrupole magnet power supplies started acting up, causing Linac beam aborts. Operators change out the chassis, but this did not solve the problem. Eventually, Linac personnel replaced a SRM. Operations reported that the Booster LLRF room AC unit failed completely. FESS personnel checked out the unit and determined that repairs would have to wait until the day shift. Operators set up a fan to blow cold outside air into the room. At 2:19 AM, the NuMI target helium flow took a large step up, which caused the beam permit to trip off. Operations contacted the NuMI machine coordinator and he decided to change over from pressure regulation to flow regulation during the day shift. Operators loaded a stacking-only timeline into the Time Line Generator (TLG) until NuMI could be fixed. Operators reported that Linac went down around 7:23 AM when a quadrupole power supply (QPS130) tripped off, followed shortly by a return of the problem from earlier in the weekend when all the tank-one quadrupoles readbacks went bad.

The Plans for Monday

The plans for today are to terminate store 8737 and allow cryo system personnel to repair flywheels on the TEV sectors A3 and F3 wet engines; this work might take up to eight hours. The Run Coordinator will also allow D0 and CDF to access their collision halls, allow MI personnel to conduct a coalescing check, and Linac personnel to conduct a I- Source flush. The Run Coordinator may also allow other work during this downtime.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/16/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory